OpenStudy (sumeer):

ax+y=1 x+ay=b does it has a solution?

OpenStudy (perl):

solve for y in the first equation. y = 1 - ax x + a(1-ax) = b x + a - a^2x = b x( 1 - a^2) = b - a x = (b-a) / ( 1 - a^2) this has a solution as long as a is not equal to 1,-1 otherwise we have division by zero which is undefined

OpenStudy (sumeer):

could you please use matrices to solve this question?

OpenStudy (perl):